find the general solutions of (x+y+2z) dz/dx+(x+y+2z)dz/dy=3z


1
Expert's answer
2022-03-30T11:35:02-0400

"(x+y+2z)\frac{\partial z}{dx}+(x+y+2z)\frac{\partial z}{dy}=3z"

The auxiliary equations is:

"\frac{dx}{x+y+2z}=\frac{dy}{x+y+2z}=\frac{dz}{3z}"

A first characteristic equation comes from

"\frac{dx}{x+y+2z}=\frac{dy}{x+y+2z}"

"x=y+C_1"

A second characteristic equation comes from

"\frac{dx+dy-\frac43dz}{2x+2y+4z-4z}=\frac{dz}{3z}"

"\frac{d(x+y)-\frac43dz}{2(x+y)}=\frac{dz}{3z}"

Set "x+y=t"

"\frac{dt-\frac43dz}{2t}=\frac{dz}{3z}"

"3zdt-4zdz=2tdz"

"3z\frac{dt}{dz}-2t=4z"

"3\frac{dt}{dz}-\frac{2}{z}t=4" (1)

Let’s solve the following equation

"3\frac{dt}{dz}-\frac{2}{z}t=0"

"3\frac{dt}{t}=2\frac{dz}{z}"

"3\ln t=2\ln z+3\ln{C(z)}"

"t=C(z)z^{\frac23}"

To find solution of the equation (1) we should think C is a function of z

"t'=C'z^{\frac23}+\frac23Cz^{-\frac13}"

"3(C'z^{\frac23}+\frac23Cz^{-\frac13})-2Cz^{-\frac13}=4"

"3C'z=4"

"C'=\frac43\frac1z"

"C=\frac43\ln |z|+C_2"

"t=(\frac43\ln |z|+C_2)z^{\frac23}"

"x+y=(\frac43\ln |z|+C_2)z^{\frac23}"

"C_2=(x+y)z^{-\frac23}-\frac43\ln |z|"

General solution of the PDE on the form of implicit equation:

"\Phi(C_1,C_2)=0".